### v2

# Do Large Language Model Understand Multi-Intent Spoken Language ?

Abstract
This study marks a significant advancement by harnessing Large Language Models (LLMs) for multi-intent spoken language understanding (SLU), proposing a unique methodology that capitalizes on the generative power of LLMs within an SLU context. Our innovative technique reconfigures entity slots specifically for LLM application in multi-intent SLU environments and introduces the concept of Sub-Intent Instruction (SII), enhancing the dissection and interpretation of intricate, multi-intent communication within varied domains. The resultant datasets, dubbed LM-MixATIS and LM-MixSNIPS, are crafted from pre-existing benchmarks. Our research illustrates that LLMs can match and potentially excel beyond the capabilities of current state-of-the-art multi-intent SLU models. It further explores LLM efficacy across various intent configurations and dataset proportions. Moreover, we introduce two pioneering metrics, Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA), to provide an in-depth analysis of LLM proficiency in this complex field.
1 Introduction
“To speak a language is to take on a world, a culture.”
— Frantz Fanon
The advancement of large language models (LLMs) has significantly accelerated progress in the field of NLP Geogle. (2023); Touvron et al. (2023). Trained on extensive datasets, these models excel in various NLP tasks, such as natural language inference, summarization, and dialog systems, through in-context learning alone Kavumba et al. (2023); Hu et al. (2022b).
Spoken language understanding (SLU) is a crucial component of task-oriented dialog systems, which aim to create a semantic frame that summarizes the user’s request. The semantic frame is specially constructed using intent detection to identify user’s intents and slot filling to extract relevant semantic constituents. Since the two sub-tasks of intent detection and slot filling are closely tied (Tur and Mori, 2011), dominant SLU systems adopt joint models to model the correlation between them (Liu and Lane, 2016; Goo et al., 2018; Qin et al., 2019).
[…]
Figure 1: An example with multi-intent detection and slot filling and “TP” denote “Transition Point”.
To deal with multi-intent scenarios, an increasing number of studies have begun to focus on modeling SLU in multi-intent settings. Xu and Sarikaya (2013) and Kim et al. (2017) first explored the multi-intent SLU. Then Qin et al. (2020a) proposed an adaptive interaction framework (AGIF) to achieve fine-grained multi-intent information integration for token-level slot filling. Qin et al. (2021b) further proposed a global-locally graph interaction network (GL-GIN) to model slot dependency and interaction between multiple intents. Recently, Huang et al. (2022) proposed a chunk-level intent detection (CLID) framework to split multi-intent into single-intent with an intent transition point, achieving a promising performance.
In the domain of multi-intent SLU, the potential of large language models has not been investigated. Although it is feasible to fine-tune LLMs for enhanced intent detection, reservations persist regarding the uncontrolled generation length of LLMs. This aspect complicates the task of accurate sequence tagging, as it may lead to error propagation and obscure the clarity of intent information. Consequently, integrating LLMs into the multi-intent SLU framework presents a notable challenge.
Addressing this, we pose several critical research questions: RQ1: Can Large Language Models effectively handle multi-intent SLU, and if so, how can they be adapted or constructed to fit within a multi-intent SLU framework? RQ2: How does the performance of Large Language Models compare with the current state-of-the-art in multi-intent SLU? RQ3: Does the size of a Large Language Model influence its performance in multi-intent SLU? RQ4: What methods can be employed to fully evaluate the performance of Large Language Models in this task?
In pursuit of answers to these questions, we propose a novel methodological approach that harnesses the strengths of LLMs for multi-intent SLU. This involves the adaptation of entity slots from conventional models to accommodate a unified framework designed for simultaneous multiple intent detection and slot filling. This approach seeks to fully harness the potential of LLMs in multi-intent task performance. Moreover, drawing from the pivotal influence of sub-sentence elements Huang et al. (2022), we introduce the Sub Intent Instruction (SII) concept. This methodology enhances LLMs’ capacity for discerning and partitioning multi-intent utterances into individual sub-intent segments, dovetailing with ’chain of thought’ reasoning Wei et al. (2022).
In summary, the contributions of this work can be succinctly outlined as follows: (1) In the model aspect, we initiate the use of LLMs within the multi-intent SLU realm, injecting innovative perspectives and addressing domain complexities methodically. (2) In the data aspect, we devise entity slots tailored for LLMs in a multi-intent SLU framework and introduce sub-intent instructions to examine sub-sentence elements’ influence. The reconstructed datasets are termed LM-MixATIS and LM-MixSNIPS. (3) In the performance aspect, the LLM approach we investigate exhibits competitive prowess, if not outright superiority, when compared to the current state-of-the-art models across several benchmarks. (4) In the metric aspect, we introduce two innovative measures: Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA), to provide a multifaceted and discerning evaluation of LLMs in the context of multi-intent SLU.
For detailed information on related work, please refer to Appendix A. Our code and datasets are available at https://github.com/SJY8460/SLM.
2 Problem Definition
2.1 Multi-Intent Detection
Given an input sequence $x=(x_{1},\ldots,x_{n})$, multi-intent detection can be defined as a multi-label classification task that outputs a sequence of intent labels $o_{I}=(o_{1}^{I},\ldots,o_{m}^{I})$, where $m$ is the number of intents in a given discourse and $n$ is the length of the discourse.
2.2 Slot Filling
Slot filling can be considered as a sequence annotation task that maps the input discourse $x$ to a slot output sequence $o_{S}=(o_{1}^{S},\ldots,o_{n}^{S})$.
Figure 2: Diagram illustrating the workflow of LLM training and inference. "Optional" denotes that sub-intent instructions are utilized only in additional explorations, not the main experiment.
3 Approach
As depicted in Figure 2, our methodology establishes a detailed framework for the integration of LLMs within the sphere of multi-intent SLU. This framework is composed of several essential components: the innovative reconfiguration of conventional slot labels to create entity slots; the introduction of Sub Intent Instruction, which bolsters the model’s grasp of complex intents; a dedicated training stage featuring supervised fine-tuning to refine model performance; the incorporation of QLoRA Quantization to streamline model scaling; and a comprehensive inference procedure to guarantee precise multi-intent SLU discernment.
3.1 Entity Slots Construction
In the Entity Slots Construction phase, we advance beyond traditional BIO sequence labeling by developing a structured entity-slot mapping specifically tailored for generative modeling with LLMs. This reconfiguration is critical for aligning with the generative capabilities of LLMs, which are more naturally inclined towards producing rich, structured outputs, as opposed to the linear nature of sequence labeling.
The process of transforming traditional BIO-annotated slots into an entity-slot mapping is mathematically articulated as follows: Consider a sequence of tokens $T=\{t_{1},t_{2},...,t_{n}\}$ alongside a corresponding sequence of BIO-annotated slots $S=\{s_{1},s_{2},...,s_{n}\}$. The transformation into entity-slot mapping is encapsulated by a function $f$:
$f(T,S)\mapsto E$ (1)
where $E=\{e_{1},e_{2},...,e_{m}\}$ represents the resultant entity-slot mapping and $m$ donates the number of entity slot within an utterance. The function $f$ processes the sequence $S$, identifying ’B-’ as the initiation of an entity and ’I-’ as its continuation. The transition to an ’O’ tag or a shift in the entity type signals the conclusion of the current entity. At this juncture, the accumulated tokens corresponding to this entity are amalgamated and incorporated into $E$ as a key-value pair. The key represents the entity type, while the value constitutes the concatenated tokens. This procedure iterates until all entities within $S$ are systematically extracted and mapped.
3.2 Sub Intent Instruction
Inspired by Huang et al. (2022)’s method of chunk-level intent detection (CLID), which segments multi-intent utterances into individual intents, we refine our model’s comprehension of intents and slots within sentences. Our approach, grounded in the practice of deconstructing multi-intent utterances into singular intents, introduces the concept of ’sub-intent’ guidance during training. By employing a chain-of-thought concept (Wei et al., 2022), our objective is to enhance the model’s capability to accurately identify and classify intents and slots.
In particular, our training dataset is enriched with sub-utterance instructions that prompt the model to concentrate on the intricate intent and slot configurations present in multi-intent utterances. This heightened level of detail during training fosters a more nuanced perception and response formulation. The training samples encompass not only the primary utterance but also a clear directive for sub-utterances associated with each respective intent.
As depicted in Figure 2, our instructional strategy mirrors the human cognitive approach of parsing complex inquiries into more manageable segments, thereby directing the model to emulate a corresponding thought process. By incorporating these sub-utterance directives, we aim to scrutinize their influence on improving LLMs’ efficacy in SLU tasks.
3.3 Supervised Fine-tuning
We utilize supervised fine-tuning to refine the generative capabilities of LLMs, ensuring they align with the detailed structures of multi-intent SLU. Initially, LLMs are trained on a broad range of data, which imparts a general understanding of language. However, to apply these models effectively in the specialized domain, supervised fine-tuning is essential. It involves the iterative adjustment of the pre-trained model’s parameters using a task-specific labeled dataset. This targeted refinement enhances the model’s accuracy and its ability to recognize and replicate the nuanced patterns unique to SLU tasks.
The fine-tuning process is defined mathematically as an optimization problem. The goal is to identify the optimal parameter set $\theta^{*}$ that minimizes the loss function $L$ over the training dataset $D$. This is expressed in the following equation:
[…]
In this context, the loss function $L$ is typically instantiated as a cross-entropy loss, which quantifies the discrepancy between the model’s predictions and the actual labels in the dataset $D$. This process is facilitated by back-propagation, which computes gradients that guide the adjustment of parameters $\theta$ towards minimizing the loss. Through this meticulous process of supervised fine-tuning, LLMs are transformed from their generalist pre-training state to become adept at handling the specific demands of multi-intent SLU.
3.4 Efficient Fine-tuning with QLoRA Quantization
In optimizing LLMs for multi-intent SLU, our approach integrates Quantized Low-Rank Adapters (QLoRA) Dettmers et al. (2023), markedly improving both efficiency and scalability. QLoRA, an advancement on the Low-Rank Adapters (LoRA) technique Hu et al. (2022a), selectively modifies a limited set of model weights. This targeted adaptation significantly reduces memory usage during training, enhancing resource management. Further, QLoRA employs quantized weights in the LoRA fine-tuning process, which substantially lowers the memory footprint.
This integration of QLoRA into our LLM fine-tuning strategy leads to a process that is not only memory-efficient but also maintains high performance, crucial for the nuanced demands of multi-intent SLU. By adopting this approach, we harness advanced capabilities of LLMs even in environments with limited computational resources, making sophisticated NLP technologies more accessible and scalable.
3.5 Inference Phase
During the inference phase for our LLMs in multi-intent SLU, we focus on processing test sentences to extract intents and entity slots. This process involves utilizing regular expressions to parse the LLM’s output, followed by a structured transformation into the standard BIO tagging format, as shown in Figure 2. The transformation process is mathematically formulated to ensure precision and clarity. Given a sentence represented by a sequence of words $T=\{t_{1},t_{2},...,t_{n}\}$, we initially tag each word with ’O’, denoting the absence of an entity. Entity slots extracted from the LLM output are represented as $E=\{e_{1},e_{2},...,e_{m}\}$. The task then is to align these slots with the words in $T$ and reassign appropriate BIO tags. The re-tagging function, denoted as $f$, operates on $T$ and $E$ to produce a sequence of BIO tags $S=\{s_{1},s_{2},...,s_{n}\}$. The function is defined as:
$f(T,E)\rightarrow S$ (3)
For each word $w_{i}$ in $S$, the function $f$ checks for alignment with entities in $E$. If $w_{i}$ marks the beginning of an entity, it is tagged as ’B-’, and if it is inside an entity, it is tagged as ’I-’. Words not part of any entity retain their ’O’ tag.
Model MixATIS MixSNIPS
Slot(F1) Intent(Acc) Overall(Acc) Slot(F1) Intent(Acc) Overall(Acc)
Stack-Propagation (Qin et al., 2019) 87.8 72.1 40.1 94.2 96.0 72.9
AGIF (Qin et al., 2020b) 86.9 72.2 39.2 93.8 95.1 72.7
GL-GIN (Qin et al., 2021b) 87.2 75.6 41.6 93.7 95.2 72.4
SDJN (Chen et al., 2022) 88.2 77.1 44.6 94.4 96.5 75.7
CLID (Huang et al., 2022) 88.2 77.5 49.0 94.3 96.6 75.0
SSRAN Cheng et al. (2023) 89.4 77.9 48.9 95.8 98.4 77.5
SDJN(BERT) 87.5 78.0 46.3 95.4 96.7 79.3
CLID(Roberta) 85.9 80.5 49.4 96.0 97.0 82.2
ChatGPT (5-shot) 64.0 54.1 14.6 62.9 83.9 12.7
Vicuna-7B-v1.5 Peng et al. (2023) 83.3 79.5 47.3 95.7 97.6 78.9
Llama-2-7B-chat Touvron et al. (2023) 86.5 82.4 51.1 95.7 96.9 78.9
Vicuna-13B-v1.5 Peng et al. (2023) 87.9 $\textbf{83.6}^{*}$ 50.8 95.9 97.5 80.7
Llama-2-13B Touvron et al. (2023) 87.9 81.0 49.5 96.7 97.8 $\textbf{83.3}^{*}$
Mistral-7B-Instruct-v0.1 Jiang et al. (2023) 88.7 80.6 $\textbf{53.4}^{*}$ 95.6 97.6 79.8
Table 1: SLU performance on MixATIS and MixSNIPS datasets. Values with * indicate that the improvement from our model is statistically significant over all baselines ($p<0.05$ under t-test).
Model intent num = 1 intent num = 2 intent num = 3
Slot(F1) Intent(Acc) Overall(Acc) Slot(F1) Intent(Acc) Overall(Acc) Slot(F1) Intent(Acc) Overall(Acc)
CLID 88.6 94.7 76.4 88.1 77.5 48.4 87.6 64.3 28.5
CLID(Roberta) 88.6 95.8 77.6 85.4 80.3 48.8 84.7 66.8 29.0
Vicuna-7B-v1.5 90.0 94.4 80.4 86.7 82.7 50.9 85.2 70.5 28.0
Llama-2-7B-chat 86.9 96.5 79.7 87.0 83.1 50.9 85.6 70.5 31.0
Mistral-7B-Instruct-v0.1 90.0 93.0 77.6 88.9 81.6 55.3 88.0 69.0 31.5
Table 2: The result comes from the dataset MixATIS. The intent num denotes the number of intents in an utterance.
4 Experiments
4.1 Datasets
Experiments were conducted on two widely-used multi-intent SLU datasets: MixATIS Hemphill et al. (1990); Qin et al. (2020a) and MixSNIPS Coucke et al. (2018); Qin et al. (2020a). The MixATIS dataset comprises 13,162 training instances, 756 validation instances, and 828 test instances, predominantly consisting of airline-related queries. Alternatively, MixSNIPS covers a broader range of domains, including restaurants, hotels, and movies, with 39,776 training instances, 2,198 validation instances, and 2,199 test instances. These datasets are designed to reflect practical scenarios, featuring utterances with 1-3 intents in a distribution ratio of 30%, 50%, and 20% respectively. Additionally, this study introduces augmented versions of these datasets by constructing entity slots and sub-intent instructions for the original data, referred to as LM-MixATIS and LM-MixSNIPS.
4.2 Experimental Settings
To optimize the training process in our studies, we applied several crucial configurations. We set the LoRA rank to 16, selected an alpha scaling parameter of 32, and established a dropout rate of 0.05. The learning rate was set to 1e-4. The training steps were determined by the corresponding dataset sizes, with 13,162 for MixATIS and 39,116 for MixSNIPS. To increase memory efficiency, we utilized the paged_adamw_8bit optimizer. For the generative configuration, the parameters were as follows: a temperature of 0.1, a top-k of 40, and a top-p of 0.75. All experiments were performed on a machine equipped with dual GeForce RTX 3090 Ti GPUs.
4.3 Baselines
In our study, we benchmark LLMs performance against a range of established baselines in the multi-intent SLU domain. These include vanilla models like Stack-Propagation (Qin et al., 2019): a stack-propagation framework to explicitly incorporate intent detection for guiding slot filling. AGIF (Qin et al., 2020b): an adaptive interaction network to achieve fine-grained multi-intent information integration, GL-GIN (Qin et al., 2021b): a local slot-aware and global intent-slot interaction graph framework to model the interaction between multiple intents and all slots within an utterance, SDJN (Chen et al., 2022): a multiple instance learning and self-distillation framework for weakly supervised multiple intent information capturing, CLID (Huang et al., 2022): a chunk-level intent detection framework for recognizing intent within a fragment of an utterance and SSRAN (Cheng et al., 2023): a transformative network built on the Transformer model, designed to reduce the complexity of multi-intent detection in SLU through scope recognition and bidirectional interaction between results of slot filling and intent detection. We also included PLM-based models like SDJN(Bert) and CLID(Roberta) which are extensions of their respective base models with pre-train language model backbones.
4.4 Large Language Model
The landscape of natural language processing has been revolutionized by the advent of open-source Large Language Models such as Llama2(Touvron et al., 2023), VicunaPeng et al. (2023), and MistralJiang et al. (2023). These models’ advanced parameterization and diverse training regimes have notably enhanced NLP task performance.
In our study, we harness state-of-the-art LLMs from these open-source initiatives, including configurations such as Vicuna-7B-v1.5 and Vicuna-13B-v1.5 (Peng et al., 2023), Llama-2-7B-chat and Llama-2-13B, and Mistral-7B-Instruct-v0.1 (Jiang et al., 2023). For comprehensive details on these LLMs, please consult Appendix A.2.
Moreover, we extend our evaluation to examine ChatGPT’s proficiency in a 5-shot learning context, paralleling the tests conducted with the LLMs outlined above. A distinctive aspect of our methodology with ChatGPT involves introducing an explicit intent and slot list to direct the model’s learning, enhancing its capability to align with specified NLP tasks.
4.5 Main Results Analysis
We evaluate performance in these areas: F1 score is used for slot filling, accuracy for intent prediction, and overall accuracy for sentence-level semantic frame parsing according to Qin et al. (2021b) and Huang et al. (2022). The overall accuracy represents the proportion of sentences in which both intent and slot are accurately predicted and it is the most important metric in multi-intent SLU.
As detailed in Table 1, our results indicate that: (1) The LLMs have manifested notable enhancements when compared with the established model CLID(Roberta). Within the single-domain MixATIS dataset, Mistral-7B-Instruct-v0.1 outstrips CLID(Roberta) by exhibiting a 2.8% increase in intent prediction accuracy, a 0.1% rise in slot filling F1 score, and a 4.0% advancement in overall sentence-level semantic frame parsing accuracy. In the multi-domain MixSNIPS dataset, Llama2-13B shows improvement over CLID(Roberta) with a 0.7% boost in intent prediction accuracy, a 0.8% increase in slot filling F1 score, and a 1.1% growth in overall accuracy. (2) When comparing the performance based on the size of the LLMs, the larger 13B models do not uniformly outperform the 7B models in the MixATIS dataset. This suggests that the model size alone may not resolve the complexity of detailed slot and intent recognition in single-domain datasets. However, the scenario is different in the multi-domain MixSNIPS dataset, where the expanded capabilities of the 13B models seem to confer an advantage. Specifically, Vicuna-13B-v1.5 improves upon its 7B counterpart by 1.8% in overall accuracy, and the Llama2-13B notably enhances performance by 4.4% in the same metric. This observation suggests that larger models can be more effective in environments with a wide range of domains, likely due to their superior ability to capture and process the diverse and intricate patterns present in the data.
5 Further Exploration
This section delves into a comprehensive analysis of the role and potential of LLMs in enhancing multi-intent SLU. Significantly, for cost-efficiency and replicability purposes, the experiments discussed within this section are uniformly conducted using 7B models.
5.1 Improvement Analysis
Large Language Models (LLMs) consistently surpass baseline models in performance across a comprehensive range of metrics within both the MixATIS and MixSNIPS datasets. To discern the performance differentials among the models, a methodical grouping was employed, predicated on the number of intents within MixATIS’s multi-intent dataset as captured in Table 2. It becomes evident that the escalation in the number of intents proportionately augments the difficulty for all models, as manifested in the diminished metrics across the board.
In a single-intent environment, the Vicuña-7B-v1.5 model demonstrates superior performance over the CLID(Roberta) model, achieving an 80.4% overall accuracy rate and marking an enhancement of 2.8%. As we venture into scenarios with two and three intents, a ubiquitous downward trend in model efficacy is noted. Yet, within these more demanding contexts, the Mistral-7B-Instruct-v0.1 model emerges as exceptional. It surpasses the CLID(Roberta) baseline by 6.5% in two-intent settings, achieving a 55.3% overall accuracy, the highest among the models under review. This distinction is sustained with three intents, where Mistral-7B-Instruct-v0.1 model attains an overall accuracy of 31.5%, topping the charts and revealing a 2.5% improvement over CLID(Roberta). The data suggests that Mistral-7B-Instruct-v0.1 possesses a significant aptitude for adeptly managing the complexities associated with multi-intent scenarios.
5.2 Sub Intent Instruction’s Efficacy in Multi-Domain SLU
Acknowledging that Large Language Models are extensively trained within diverse domains, our investigation extends to the Sub Intent Instruction (SII) and its role in bolstering LLMs’ proficiency for discerning and segmenting multi-intent utterances into discrete sub-intent components within the multi-domain MixSNIPS dataset. This enhancement ostensibly refines the overall multi-intent Spoken Language Understanding (SLU) capability—a claim substantiated by the data in Table 3.
Upon the integration of SII, although it does not improve significantly on Intent(Acc) and Slot(f1), a substantial increase in Overall Accuracy has been observed across all models, underlining SII’s pivotal role in boosting joint intent and slot recognition efficacy. For instance, Vicuna-7B-V1.5’s Overall(Acc) witnessed a significant rise of 3.5%, climbing from 78.9% to 82.4%. A similar enhancement was seen with Llama-2-7B-Chat, which recorded a robust improvement of 4.6%, reaching an Overall(Acc) of 83.5%. Mistral-7B-Instruct-v0.1 also experienced a notable gain of 4.1% in Overall(Acc), achieving 83.9%.
These pronounced increments in Overall(Acc) with the integration of SII emphasize its pivotal significance in ameliorating multi-intent SLU performance, necessitating accurate identification of the intricate links between various intents and their corresponding slots.
Model Slot(F1) Intent(Acc) Overall(Acc)
Vicuna-7B-v1.5 95.7 97.6 78.9
+SII 96.2 (↑0.5) 97.1 (↓0.5) 82.4 (↑3.5)
Llama-2-7B-chat 95.7 96.9 78.9
+SII 96.5 (↑0.8) 97.6 (↑0.7) 83.5 (↑4.6)
Mistral-7B-Instruct-v0.1 95.6 97.6 79.8
+SII 96.7 (↑1.1) 97.5 (↓0.1) 83.9 (↑4.1)
Table 3: Further experiments on the MixSNIPS dataset showing the impact of the sub intent instruction (SII). The brackets indicate the change in performance metrics with the SII approach compared to the original model.
Figure 3: Performance comparison of models on MixATIS and MixSNIPS datasets. The figures depict Semantic Accuracy, Slot F1, and Intent Accuracy across different data ratios for Mistral, Vicuna, and Llama models (7B).
5.3 Correlation between Training Data Volume and Model Efficacy
As illustrated in Figure 3, we probes the effect of varying volumes of training data on model proficiency, with a focus on three critical performance indicators: Semantic Accuracy, Slot F1 Score, and Intent Accuracy. Through progressive modulation of training data ratios at intervals of 0.2, 0.4, 0.6, 0.8, and 1.0. There appears to be a consistent and positive correlation between the data ratio and the amelioration of performance metrics across both the MixATIS and MixSNIPS datasets. This trend substantiates the hypothesis that an increase in dataset size can significantly bolster the model’s training efficacy, thereby refining accuracy and broadening the model’s generalization abilities. Upon closer inspection, the MixATIS and MixSNIPS datasets indicate that the Mistral-7B-Instruct-v0.1 model, starting at a less optimal performance with a data ratio of 0.2, demonstrates notable advancements as the dataset size increases. For instance, within the MixSNIPS dataset, there is an impressive ascent in Semantic Accuracy from 51.16% to 79.76%, Slot F1 from 76.18% to 95.64%, and Intent Accuracy from 72.99% to 97.64%. These substantial improvements underscore the model’s ability to enhance its learning and semantic processing capabilities with access to a larger corpus of data. Meanwhile, in the MixATIS dataset, Mistral-7B-Instruct-v0.1 achieves promising results in Semantic Accuracy and Slot F1. However, it appears to have compromised some aspect of Intent Accuracy in comparison to its counterparts, Llama2 and Vicuna.
6 Constructing Metrics for Evaluating LLMs in SLU
To enhance the evaluation of LLMs in the domain of multi-intent SLU, we propose two novel metrics for a more nuanced analysis: Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA).
Entity Slot Accuracy (ESA): ESA measures the model’s ability to precisely identify entity slots within utterances, which are crucial for understanding the specific semantic elements of the spoken input.
Combined Semantic Accuracy (CSA): CSA amalgamates intent accuracy with ESA to yield a comprehensive gauge of a model’s capability in discerning the overall semantic framework of utterances. The metric is mathematically represented as follows:
$CSA=\text{Intent Accuracy}\times\text{ESA}$ (4)
Upon analyzing the performance of various Language Learning Models (LLMs) on the MixATIS and MixSNIPS datasets, the data reveals certain performance trends. The Mistral-7B-Instruct-v0.1 model outshines others in the MixATIS dataset with an ESA of 60.6%, and a CSA of 48.8%. In contrast, the Llama-2-13B model demonstrates a remarkable aptitude in the MixSNIPS dataset, achieving the highest ESA at 83.6% and the highest CSA at 81.7%. These scores reflect the model’s proficiency in handling the complexity of the dataset’s queries. It is worth noting that the ESA and CSA metrics provide a refined assessment of model performance by exclusively considering the value of entity slots, without taking into account the "O" tag typically used to denote tokens that do not correspond to any entity slot in the dataset’s utterances. This focus results in a more rigorous evaluation of an LLM’s ability to recognize and accurately predict the relevant information within a given input.
Model MixATIS MixSNIPS
ESA CSA ESA CSA
Vicuna-7B-v1.5 53.3 43.5 79.1 77.2
Llama-2-7B-chat 53.1 43.8 79.3 76.8
Vicuna-13B-v1.5 56.2 46.9 80.9 78.9
Llama-2-13B 59.5 48.3 83.6 81.7
Mistral-7B-Instruct-v0.1 60.6 48.8 80.2 78.3
Table 4: ESA and CSA performance of various models on the MixATIS and MixSNIPS datasets.
7 Conclusion
In this study, we have integrated LLMs into the multi-intent SLU framework, marking a leap forward in four critical areas: (1) In terms of modeling, we pioneered the application of LLMs for multi-intent SLU, offering fresh insights and systematically tackling the inherent complexities of the domain. (2) From a data perspective, we crafted entity slots optimized for LLMs within the multi-intent SLU context and established sub-intent instructions to probe the impact of sub-sentence components. The datasets refined for this purpose are named LM-MixATIS and LM-MixSNIPS. (3) Regarding performance, our investigation reveals that the LLM-based approach demonstrates impressive competitiveness, and in many instances, surpasses the capabilities of the existing state-of-the-art models in various benchmarks. (4) In the domain of metrics, we have introduced two novel measures: Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA), to facilitate a comprehensive and nuanced assessment of LLMs within multi-intent SLU. Through these advancements, we aspire to catalyze greater interest and exploration into the potential of LLMs to enhance multi-intent SLU systems.
8 Limitations
In this paper, we openly acknowledge several limitations of our approach: (1) Quantization Impact on Performance: Our implementation leverages QLoRA quantization to enhance training efficiency and reduce memory consumption. However, this approach might not reach the full efficacy of complete-parameter fine-tuning. The inherent trade-off in quantization is efficiency at the potential expense of some performance metrics. While it streamlines the process, it may inadvertently overlook subtle nuances that full-parameter models can capture. (2) Anticipated Enhancements through Data Selection and Prompt Refinement: Our current study does not encompass the advanced techniques of meticulous data selection or sophisticated prompt engineering. Recognizing this as a limitation, our future work will aim to incorporate these critical strategies. By deliberately curating datasets and enhancing prompt design in subsequent research, we intend to further refine the model’s performance, especially in areas where quantization may have compromised subtleties in linguistic understanding. These anticipated improvements will strive to strike a balance between efficiency and the depth of language comprehension, presenting a more comprehensive solution within the domain of Large Language Models.
Appendix A Related Work
A.1 Intent Detection and Slot Filling
Intent detection and slot filling are often interrelated, giving rise to the development of integrated models that facilitate interaction between intent and slots. In recent years, techniques such as joint learning, which consider the strong correlation between intent and slots, have yielded outstanding results. Certain approaches to joint slot filling and intent detection involve the sharing of parameters (Liu and Lane, 2016; Wang et al., 2018; Zhang and Wang, 2016). The association between intent detection and slot filling can be modeled through unidirectional interaction or bidirectional-flow interaction (Qin et al., 2021c).
Unidirectional interaction approaches (Goo et al., 2018; Li et al., 2018; Qin et al., 2019) primarily focus on the flow from intent to slot. Gating mechanisms have been employed as specialized functions to guide slot filling (Goo et al., 2018; Li et al., 2018). Qin et al. (2019) proposed a token-level intent detection model to mitigate error propagation.
Bidirectional-flow interaction models (E et al., 2019; Zhang et al., 2019; Liu et al., 2019; Qin et al., 2021a) consider the mutual impact between intent detection and slot filling. E et al. (2019) enhanced intent detection and slot filling bidirectionally through iteration mechanisms.
More recently, Chen et al. (2022) introduced a Self-distillation Joint SLU model, leveraging multi-task learning. They also treated multiple intent detection as a weakly supervised challenge, employing multiple instance learning (MIL). Huang et al. (2022) proposed a chunk-level intent detection framework, including an auxiliary task to identify intent transition points within utterances, thereby enhancing the recognition of multiple intents. And Cheng et al. (2023) proposed a transformative network built on the Transformer model, designed to reduce the complexity of multi-intent detection in SLU through scope recognition and bidirectional interaction between results of slot filling and intent detection.
A.2 Open Source Large Language Model
Advancements in natural language processing (NLP) have been significantly driven by the development of open-source Large Language Models (LLMs) like Llama2(Touvron et al., 2023), VicunaPeng et al. (2023), and MistralJiang et al. (2023). These models, with their extensive parameterization and training on diverse datasets, have expanded the capabilities in various NLP tasks.
Llama2, a creation of Meta AI, offers a range of models with parameters varying from 7 billion to 70 billion. It excels particularly in dialogue-based applications, utilizing auto-regressive language optimized transformers and advanced techniques such as supervised fine-tuning and reinforcement learning. This approach enhances Llama2’s performance, especially in dialogue generation and understandingTouvron et al. (2023). Vicuna, another notable open-source LLM, is specially designed for chatbot applications. It is trained on datasets that mirror real-world interactions, such as those from ShareGPT.com. Vicuna stands out for its transparent approach to training and its architectural innovations, offering various versions like Vicuna-13B to cater to different application needsPeng et al. (2023). Mistral 7B, developed by Mistral AI, is geared towards both English language processing and coding tasks. This 7.3 billion parameter model is distinguished for its superior performance over Llama 2 13B and Llama 1 34B in multiple benchmarks. The efficiency of Mistral 7B can be attributed to its use of grouped-query attention, which balances resource efficiency with output qualityJiang et al. (2023).
Chain of Thought Wei et al. (2022) is yet another LLM gaining attention for its capabilities in language generation. Together, these models, including Llama2, Vicuna, Mistral, and Chain of Thought, contribute to a diverse landscape of LLMs. They not only enhance NLP tasks but also have the potential to revolutionize content creation and search engine utilization.