### v3

# Do Large Language Model Understand Multi-Intent Spoken Language ?

Abstract
This research signifies a considerable breakthrough in leveraging Large Language Models (LLMs) for multi-intent spoken language understanding (SLU). Our approach re-imagines the use of entity slots in multi-intent SLU applications, making the most of the generative potential of LLMs within the SLU landscape, leading to the development of the EN-LLM series. Furthermore, we introduce the concept of Sub-Intent Instruction (SII) to amplify the analysis and interpretation of complex, multi-intent communications, which further supports the creation of the ENSI-LLM models series. Our novel datasets, identified as LM-MixATIS and LM-MixSNIPS, are synthesized from existing benchmarks. The study evidences that LLMs may match or even surpass the performance of the current best multi-intent SLU models. We also scrutinize the performance of LLMs across a spectrum of intent configurations and dataset distributions. On top of this, we present two revolutionary metrics - Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA) - to facilitate a detailed assessment of LLM competence in this multifaceted field." Our code and datasets are available at https://github.com/SJY8460/SLM.
1 Introduction
“To speak a language is to take on a world, a culture.”
— Frantz Fanon
The advancement of large language models (LLMs) has significantly accelerated progress in the field of NLP Geogle. (2023); Touvron et al. (2023). Trained on extensive datasets, these models excel in various NLP tasks, such as natural language inference, summarization, and dialog systems, through in-context learning alone Kavumba et al. (2023); Hu et al. (2022b).
Spoken language understanding (SLU) is a crucial component of task-oriented dialog systems, which aim to create a semantic frame that summarizes the user’s request. The semantic frame is specially constructed using intent detection to identify user’s intents and slot filling to extract relevant semantic constituents. Since the two sub-tasks of intent detection and slot filling are closely tied (Tur and Mori, 2011), dominant SLU systems adopt joint models to model the correlation between them (Goo et al., 2018; Qin et al., 2019).
Figure 1: An example with Multi-Intent SLU, where B-WT donates B-Weather, B-LOC donates B-Location and “TP” denote “Transition Point”.
[…]
Recently, Huang et al. (2022) proposed a chunk-level intent detection (CLID) framework to split multi-intent into single-intent with an intent transition point. Furthermore, Yin et al. (2024) develop an united multi-view intent-slot interaction framework(Uni-MIS), achieving promising performance.
In the domain of multi-intent SLU, the potential of large language models has not been investigated. Although it is feasible to fine-tune LLMs for enhanced intent detection, reservations persist regarding the uncontrolled generation length of LLMs. This aspect complicates the task of accurate sequence tagging, as it may lead to error propagation and obscure the clarity of intent information. Consequently, integrating LLMs into the multi-intent SLU framework presents a notable challenge.
Addressing this, we pose several critical research questions: RQ1: Can Large Language Models effectively handle multi-intent SLU, and if so, how can they be adapted or constructed to fit within a multi-intent SLU framework? RQ2: How does the performance of Large Language Models compare with the current state-of-the-art in multi-intent SLU? RQ3: What methods can be employed to fully evaluate the performance of Large Language Models in this task?
To address these queries, we introduce a novel method, leveraging the capabilities of LLMs for multi-intent SLU. This method adapts the entity slots from traditional models, thus offering a unified framework designed for simultaneous multiple intent detection and slot filling. As depicted in Fig 1, the task of token-level slot labeling, often termed as slot filling, can be further refined into an entity-level slot detection. This enables LLMs to focus on identifying pertinent slot labels without the necessity to label all slot tokens within a sentence, thereby mitigating the problem of uncontrolled length generation. Furthermore, driven by "Chain of Thought" reasoning Wei et al. (2022), we usher in the concept of Sub Intent Instruction (SII). This strategy enhances LLMs’ proficiency in distinguishing and segregating multi-intent utterances into individual sub-intent segments. For a more comprehensive evaluation of our proposed method, we introduce two innovative metrics: Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA). These metrics offer a thorough and insightful assessment of LLMs in the context of multi-intent SLU at the entity-level.
In summary, the contributions of this work can be succinctly outlined as follows: (1) In the model aspect, we initiate the use of LLMs within the multi-intent SLU realm, injecting innovative perspectives and addressing domain complexities methodically. (2) In the data aspect, we devise entity slots tailored for LLMs in a multi-intent SLU framework and introduce sub-intent instructions to examine sub-sentence elements’ influence. The reconstructed datasets are termed LM-MixATIS and LM-MixSNIPS. (3) In the performance aspect, the LLM approach we investigate exhibits competitive prowess, if not outright superiority, when compared to the current state-of-the-art models across several benchmarks. (4) In the metric aspect, we introduce two innovative measures: Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA), to provide a multifaceted and discerning evaluation of LLMs in the context of multi-intent SLU.
2 Problem Definition
2.1 Multi-Intent Detection
Given an input sequence $x=(x_{1},\ldots,x_{n})$, multi-intent detection can be defined as a multi-label classification task that outputs a sequence of intent labels $o_{I}=(o_{1}^{I},\ldots,o_{m}^{I})$, where $m$ is the number of intents in a given discourse and $n$ is the length of the discourse.
2.2 Slot Filling
Slot filling can be considered as a sequence annotation task that maps the input discourse $x$ to a slot output sequence $o_{S}=(o_{1}^{S},\ldots,o_{n}^{S})$.
Figure 2: Diagram illustrating the workflow of LLM training and inference. Sub-Intent instruction is only used in ENSI models.
3 Approach
As depicted in Figure 2, our methodology establishes a detailed framework for the integration of LLMs within the sphere of multi-intent SLU. This framework is composed of several essential components: the innovative reconfiguration of conventional slot labels to create entity slots; the introduction of Sub Intent Instruction, which bolsters the model’s grasp of complex intents; and a comprehensive inference procedure to guarantee precise multi-intent SLU discernment.
3.1 Entity Slots Construction
In the Entity Slots Construction phase, we advance beyond traditional BIO sequence labeling by developing a structured entity-slot mapping specifically tailored for generative modeling with LLMs. This reconfiguration is critical for aligning with the generative capabilities of LLMs, which are more naturally inclined towards producing rich, structured outputs, as opposed to the linear nature of sequence labeling.
The process of transforming traditional BIO-annotated slots into an entity-slot mapping is mathematically articulated as follows: Consider a sequence of tokens $T=\{t_{1},t_{2},...,t_{n}\}$ alongside a corresponding sequence of BIO-annotated slots $S=\{s_{1},s_{2},...,s_{n}\}$. The transformation into entity-slot mapping is encapsulated by a function $c$:
$c(T,S)\mapsto E$ (1)
where $E=\{e_{1},e_{2},...,e_{m}\}$ represents the resultant entity-slot mapping and $m$ donates the number of entity slot within an utterance. The function $c$ processes the sequence $S$, identifying ’B-’ as the initiation of an entity and ’I-’ as its continuation. The transition to an ’O’ tag or a shift in the entity type signals the conclusion of the current entity. At this juncture, the accumulated tokens corresponding to this entity are amalgamated and incorporated into $E$ as a key-value pair. The key represents the entity type, while the value constitutes the concatenated tokens. This procedure iterates until all entities within $S$ are systematically extracted and mapped.
3.2 Sub-Intent Instruction
Inspired by the process of chunk-level intent detection (CLID)(Huang et al., 2022), which breaks down multi-intent utterances into singular intentions, and guided by the "Chain of Thought" reasoning Wei et al. (2022), we refine our model’s discernment of intents within sentences. Our methodology is grounded in the systematic dissection of multi-intent statements into distinct intents. For an instance dealing with two intents, the Sub-Intent Instruction could be represented as {"Sub-Intent-1": Sub-Utterance-1; "Sub-Intent-2": Sub-Utterance-2}. This enhanced granularity during the training phase fosters a more refined understanding and formulation of responses. Hence, the training samples not only include the primary utterance but also precise directives for sub-utterances correlated with each individual intent, facilitating the opportunity for improved comprehension and professionalism..
3.3 Efficient Fine-tuning with LoRA
In an effort to generate more cost-efficient experiments, we leverage advancements such as Low-Rank Adapters (LoRA) Hu et al. (2022a). Our attention shifts towards constraining the modifications in an existing weight matrix $W_{0}\in R^{d\times k}$. Through a low-rank decomposition strategy, these alterations are maneuvered towards $W_{0}+\Delta W=W_{0}+BA$, where $B\in R^{d\times r}$,$A\in R^{r\times k}$, and and the rank $r\ll min(d,k)$. As the training process progresses, the consistency of $W_{0}$ is maintained, with the flexible parameters being assigned to matrices $A$ and $B$ exclusively. This methodology culminates in an adapted forward pass, defined as:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (2)
Maintaining true to LoRA’s approach, the initialization stage employs a random Gaussian for A and zero for B, leading to $\Delta W=BA$ starting from zero at inception. The selective modification of a restricted set of model weights keeps the core structure intact, whilst providing ample room for focused enhancements. To further conserve computational resources for larger scale models, we also incorporate quantization techniques (Dettmers et al., 2023).
3.4 Extract and Recovery Inference
During the inference phase, LLMs in multi-intent SLU are primarily engaged in processing test sentences for the extraction of intents and entity slots. Notably, this process leverages the utilization of regular expressions to parse the first appearing intent and entity slot information from the LLM’s output, which is necessary due to the unpredictable generation length of the LLMs. This information is subsequently obligingly transformed in a structured manner into the standard BIO tagging format, a process illustrated in Figure 2.
The transformation process is mathematically formulated to ensure precision and clarity. Given a sentence represented by a sequence of words $T=\{t_{1},t_{2},...,t_{n}\}$, we initially tag each word with ’O’, denoting the absence of an entity. Entity slots extracted from the LLM output are represented as $E=\{e_{1},e_{2},...,e_{m}\}$. The task then is to align these slots with the words in $T$ and reassign appropriate BIO tags. The re-tagging function, denoted as $r$, operates on $T$ and $E$ to produce a sequence of BIO tags $S=\{s_{1},s_{2},...,s_{n}\}$. The function is defined as:
$r(T,E)\rightarrow S$ (3)
For each word $w_{i}$ in $S$, the function $f$ checks for alignment with entities in $E$. If $w_{i}$ marks the beginning of an entity, it is tagged as ’B-’, and if it is inside an entity, it is tagged as ’I-’. Words not part of any entity retain their ’O’ tag.
Model MixATIS MixSNIPS
Slot(F1) Intent(Acc) Overall(Acc) Slot(F1) Intent(Acc) Overall(Acc)
AGIF (Qin et al., 2020b) 86.9 72.2 39.2 93.8 95.1 72.7
GL-GIN (Qin et al., 2021b) 87.2 75.6 41.6 93.7 95.2 72.4
SDJN (Chen et al., 2022) 88.2 77.1 44.6 94.4 96.5 75.7
CLID (Huang et al., 2022) 88.2 77.5 49.0 94.3 96.6 75.0
SSRAN Cheng et al. (2023) 89.4 77.9 48.9 95.8 98.4 77.5
SDJN${}_{\text{(Bert)}}$ (Chen et al., 2022) 87.5 78.0 46.3 95.4 96.7 79.3
CLID${}_{\text{(Roberta)}}$ (Huang et al., 2022) 85.9 80.5 49.4 96.0 97.0 82.2
Uni-MIS${}_{\text{(Roberta)}}$ Yin et al. (2024) 88.3 78.5 52.5 96.4 97.2 83.4
EN-Vicuna (Ours) 83.3 79.5 47.3 95.7 97.6 78.9
EN-Llama-2 (Ours) 86.5 82.4 51.1 95.7 96.9 78.9
EN-Mistral (Ours) 88.7 80.6 $\textbf{53.4}^{*}$ 95.6 97.6 79.8
ENSI-Vicuna (Ours) 86.0 80.6 49.4 96.2 97.1 82.4
ENSI-Llama-2 (Ours) 84.7 79.3 46.6 96.5 97.6 83.5
ENSI-Mistral (Ours) 86.4 78.0 49.8 96.7 97.5 83.9
Table 1: Multi-Intent SLU performance on MixATIS and MixSNIPS datasets. Values with * indicate that the improvement from our model is statistically significant over all baselines ($p<0.05$ under t-test).
4 Experiments
4.1 Datasets
Experiments were conducted on two widely-used multi-intent SLU datasets: MixATIS Hemphill et al. (1990); Qin et al. (2020a) and MixSNIPS Coucke et al. (2018); Qin et al. (2020a). The MixATIS dataset comprises 13,162 training instances, 756 validation instances, and 828 test instances, predominantly consisting of airline-related queries. Alternatively, MixSNIPS covers a broader range of domains, including restaurants, hotels, and movies, with 39,776 training instances, 2,198 validation instances, and 2,199 test instances. These datasets are designed to reflect practical scenarios, featuring utterances with 1-3 intents in a distribution ratio of 30%, 50%, and 20% respectively. Additionally, this study introduces augmented versions of these datasets by constructing entity slots and sub-intent instructions for the original data, referred to as LM-MixATIS and LM-MixSNIPS.
4.2 Experimental Settings
To optimize the training process in our studies, we applied several crucial configurations. We set the LoRA rank to 16, selected an alpha scaling parameter of 32, and established a dropout rate of 0.05. The learning rate was set to 1e-4. The training steps were determined by the corresponding dataset sizes, with 13,162 for MixATIS and 39,116 for MixSNIPS. To increase memory efficiency, we utilized the paged_adamw_8bit optimizer. For the generative configuration, the parameters were as follows: a temperature of 0.1, a top-k of 40, and a top-p of 0.75. All experiments were performed on a machine equipped with dual GeForce RTX 3090 Ti GPUs.
4.3 Baselines
In our study, we benchmark LLMs performance against a range of established baselines in the multi-intent SLU domain. These include vanilla models like Stack-Propagation (Qin et al., 2019): a stack-propagation framework to explicitly incorporate intent detection for guiding slot filling. AGIF (Qin et al., 2020b): an adaptive interaction network to achieve fine-grained multi-intent information integration, GL-GIN (Qin et al., 2021b): a local slot-aware and global intent-slot interaction graph framework to model the interaction between multiple intents and all slots within an utterance, SDJN (Chen et al., 2022): a multiple instance learning and self-distillation framework for weakly supervised multiple intent information capturing, CLID (Huang et al., 2022): a chunk-level intent detection framework for recognizing intent within a fragment of an utterance and SSRAN (Cheng et al., 2023): a transformative network built on the Transformer model, designed to reduce the complexity of multi-intent detection in SLU through scope recognition and bidirectional interaction between results of slot filling and intent detection. We also included PLM-based models such as Uni-MIS (Yin et al., 2024): a unified multi-intent slu framework via multi-view intent-slot interaction. Additionally, SDJN(Bert) and CLID(Roberta) extend their respective base models by incorporating pre-trained language model backbones.
4.4 Large Language Model
In our study, we harness state-of-the-art LLMs from these open-source initiatives, including configurations such as Vicuna-7B-v1.5(Peng et al., 2023), Llama-2-7B-chat and Mistral-7B-Instruct-v0.1 (Jiang et al., 2023). We also report the performance of larger scale open source language models and ChatGPT(few-shot) in appendix A.
4.5 Detailed Analysis of Main Results
For a robust evaluation of performance criteria, various metrics were adopted, namely: the F1 score for slot filling, accuracy for intent prediction, and overall accuracy for sentence-level semantic frame parsing, in alignment with the methodologies proposed by Qin et al. (2021b) and Huang et al. (2022). The paramount metric in multi-intent SLU is the overall accuracy, representing the proportion of sentences where both intent and slot are precisely predicted.
In this study, the term ‘EN-LLM model’ refers to LLM adapted with an entity slot, whereas ’ENSI-LLM’ models incorporate additional instructions for sub-intent. As manifested in Table 1, our findings underscore several observations:
(1) LLMs have demonstrated significant advancements compared to the prevalent model Uni-MIS. For instance, in the single-domain MixATIS dataset, EN-Mistral surpasses Uni-MIS by demonstrating a 0.4% enhancement in intent prediction accuracy, accompanied by a 2.1% improvement in slot filling F1 score, and a 0.9% uplift in overall sentence-level semantic frame parsing accuracy. Similarly, within the multi-domain MixSNIPS dataset, ENSI-Mistral outperforms Uni-MIS, reflecting a 1.1% increase in intent prediction accuracy and a 0.5% surge in overall accuracy. Our LLM models’ performance has also proven impressive when benchmarked against other formidable baseline models.
(2) With the integration of SII in the MixSNIPS datasets, significant growth is observed in overall accuracy rates. ENSI-Vicuna’s Overall (Acc) records a substantial increase of 3.5%, elevating from 78.9% to 82.4%. ENSI-Llama-2 follows a similar trajectory with a robust enhancement of 4.6%, achieving an Overall(Acc) of 83.5%. ENSI-Mistral benefits from a critical gain of 4.1% in Overall(Acc), culminating at 83.9%. However, a contrast is seen within the MixATIS dataset, where a slight drop is observed. This suggests that while SII proves beneficial for multi-domain tasks, it is less suited for single-domain tasks due to knowledge complexity. An LLM, when faced with an array of similar information, can lose accuracy in the prediction of intents and slots.
Model intent num = 1 intent num = 2 intent num = 3
Slot(F1) Intent(Acc) Overall(Acc) Slot(F1) Intent(Acc) Overall(Acc) Slot(F1) Intent(Acc) Overall(Acc)
CLID(Roberta) 88.6 95.8 77.6 85.4 80.3 48.8 84.7 66.8 29.0
Uni-MIS 89.2 95.1 78.6 87.6 78.3 50.5 86.7 66.7 31.7
EN-Vicuna 90.0 94.4 80.4 86.7 82.7 50.9 85.2 70.5 28.0
EN-Llama-2 86.9 96.5 79.7 87.0 83.1 50.9 85.6 70.5 31.0
EN-Mistral 90.0 93.0 77.6 88.9 81.6 55.3 88.0 69.0 31.5
Table 2: The result comes from the dataset MixATIS. The intent num denotes the number of intents in an utterance.
Model intent num = 1 intent num = 2 intent num = 3
Slot(F1) Intent(Acc) Overall(Acc) Slot(F1) Intent(Acc) Overall(Acc) Slot(F1) Intent(Acc) Overall(Acc)
EN-Vicuna 95.1 97.6 88.4 96.0 97.9 79.9 95.4 96.8 67.6
EN-Llama 95.5 97.3 89.3 96.0 97.1 80.6 95.0 95.8 65.0
EN-Mistral 95.8 97.8 90.2 96.1 98.1 81.3 94.8 96.2 65.8
Table 3: The result comes from the dataset MixSNIPS. The intent num denotes the number of intents in an utterance.
5 Further Exploration
This section centres around an in-depth exploration of the LLM’s potential in advancing multi-intent SLU. Emphasizing consistency, we confine our experimental studies to the use of the EN-LLM series without incorporating human-written sub-intent instruction.
5.1 Influence of Different Intent Number
Within the ambit of both the MixATIS and MixSNIPS datasets, LLMs have exhibited a substantial competitive prowess over the baseline models, spanning a diverse range of performance metrics. In order to dissect the subtleties of these performance divergences, a systematic classification was adopted on the basis of the number of intents present in the MixATIS and MixSNIPS multi-intent datasets, as duly outlined in Table 2 and Table 3. Nonetheless, for the convenience of comparison with prior studies such as (Huang et al., 2022; Yin et al., 2024), this section primarily places its focus on the MixATIS dataset, though analogous conclusions can be drawn from the MixSNIPS dataset.
It becomes apparent that increased numbers of intents correspondingly escalate the challenge for all models, resulting in overall decreased performance metrics. In the context of a single-intent setting, the EN-Vicuña model exhibits a performance superiority over the Uni-MIS model by achieving an 80.4% overall accuracy rate, reflecting an advancement of 1.8%. And the EN-Mistral model unravels a noteworthy performance by outshining the Uni-MIS baseline by 4.8% in two-intent scenarios, thereby securing an overall accuracy of 55.3%. This leading position continues even into the three-intent setting, with the EN-Mistral model reaching an overall accuracy of 31.5%, which marginally underperforms Uni-MIS by only 0.2%.
Drawing from such observations, it suggests that LLMs exhibit a substantial capacity for proficiently handling complexities associated with a small number of intents. However, as the quantity of intents escalates, most models begin to display regressing performance, potentially attributable to the amplified error propagation in LLM’s generation process as a result of its autoregressive architecture.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Performance comparison of models on MixATIS and MixSNIPS datasets. The figures depict Semantic Accuracy, Slot F1, and Intent Accuracy across different data ratios for EN-Mistral, EN-Vicuna, and EN-Llama-2 models.
5.2 Influence of Training Data Ratio
As illustrated in Figure 3, we probes the effect of varying volumes of training data on model proficiency, with a focus on three critical performance indicators: Semantic Accuracy, Slot F1 Score, and Intent Accuracy. Through progressive modulation of training data ratios at intervals of 0.2, 0.4, 0.6, 0.8, and 1.0. There appears to be a consistent and positive correlation between the data ratio and the amelioration of performance metrics across both the MixATIS and MixSNIPS datasets. This trend substantiates the hypothesis that an increase in dataset size can significantly bolster the model’s training efficacy, thereby refining accuracy and broadening the model’s generalization abilities. Upon closer inspection, the MixATIS and MixSNIPS datasets indicate that the EN-Mistral model, starting at a less optimal performance with a data ratio of 0.2, demonstrates notable advancements as the dataset size increases.
6 Introduction of ESA and CSA Metrics
To enhance the evaluation of LLMs in the domain of multi-intent SLU, we propose two novel metrics for a more nuanced analysis: Entity Slot Accuracy and Combined Semantic Accuracy.
Entity Slot Accuracy (ESA): ESA is a key metric in the evaluation of LLM performance for multi-intent SLU tasks. Mathematically, ESA is determined by the ratio of correctly identified entity slots, denoted as $E_{c}$, to the total number of slots forecasted, denoted as $E_{t}$. The quantification can be articulated by:
$ESA=\frac{|E_{c}|}{|E_{t}|}\times 100\%$ (4)
Here, ||$\cdot$|| ⋅ | signifies the cardinality or the count of elements in a set. This concise representation of ESA hence encapsulates the model’s precision in parsing the semantic features of an utterance.
Combined Semantic Accuracy (CSA): CSA aims to provide an encompassing evaluation of a model’s adeptness at understanding both the intent and entity slots in an utterance. Denoting Intent Accuracy as $I_{a}$ and Entity Slot Accuracy as $E_{s}$, CSA is then mathematically represented as:
$CSA=I_{a}\times E_{s}$ (5)
This equation essentially encapsulates the model’s holistic accuracy in correctly identifying both the overall intent and the specific slots within a given utterance.
Upon analyzing the performance of various LLMs on the LM-MixATIS and LM-MixSNIPS datasets, the data reveals certain performance trends. The EN-Mistral model outshines others in the MixATIS dataset with an ESA of 60.6%, and a CSA of 48.8%. In the MixSNIPS dataset, EN-Mistral model achieves the highest ESA at 84.2% and the highest CSA at 82.2%. These scores reflect the model’s proficiency in handling the complexity of the dataset’s queries.
Notably, ESA and CSA provide a nuanced performance assessment, particularly focusing on the value of entity slots, and excluding the ’O’ tag, typically denoting tokens not corresponding to any entity slot in the utterance. This stringent focus endorses a more thorough evaluation of an LLM’s capabilities in discerning and accurately predicting relevant utterance information.
Model MixATIS MixSNIPS
ESA CSA ESA CSA
EN-Vicuna 53.3 43.5 79.1 77.2
EN-Llama-2 53.1 43.8 79.3 76.8
EN-Mistral 60.6 48.8 80.2 78.3
ENSI-Vicuna 56.3 45.3 82.9 80.6
ENSI-Llama-2 50.0 39.7 83.9 81.9
ENSI-Mistral 53.1 41.5 84.2 82.2
Table 4: ESA and CSA performance of various models on the MixATIS and MixSNIPS datasets.
7 Related Work
7.1 Intent Detection and Slot Filling
The inherent interconnectedness of intent detection and slot filling has spurred the development of unified models that foster mutual interaction between the two elements. Joint learning techniques, acknowledging the potent correlation between intents and slots, have proven particularly efficacious in recent years. Certain methodologies facilitating simultaneous slot filling and intent detection employ shared parameters (Liu and Lane, 2016; Wang et al., 2018; Zhang and Wang, 2016), while others model the relationship between the two via either unidirectional interaction or bidirectional-flow interaction (Qin et al., 2021c).
Models adopting unidirectional interaction, such as those by (Goo et al., 2018; Li et al., 2018; Qin et al., 2019), primarily emphasize the flow from intent to slot. Gating mechanisms, functioning as specialized guiding forces for slot filling, have seen extensive use (Goo et al., 2018; Li et al., 2018). Qin et al. (2019) put forth a token-level intent detection model to curtail error propagation.
Bidirectional-flow interaction models (E et al., 2019; Zhang et al., 2019; Liu et al., 2019; Qin et al., 2021a), on the other hand, examine the reciprocal influence of intent detection and slot filling. E et al. (2019) utilized iterative mechanisms to enhance intent detection and slot filling in both directions.
Fine-grained intent detection and intent-slot interaction models have also seen remarkable advancements. Chen et al. (2022) developed a Self-distillation Joint SLU model exploitating multi-task learning, and treated multiple intent detection as a weakly-supervised problem solved through Multiple Instance Learning (MIL). Similarly, Huang et al. (2022) introduced a chunk-level intent detection framework that employs an auxiliary task to pinpoint intent transition points within utterances, thereby augmenting the recognition of multiple intents. Furthermore, Cheng et al. (2023) proposed a transformative network rooted in the Transformer model, designed to diminish the complexity of multi-intent detection in SLU. Recently, Yin et al. (2024) further develop an united multi-view intent-slot interaction framework(Uni-MIS), archiving promising performance.
7.2 Open Source Large Language Model
The landscape of Natural Language Processing (NLP) has been notably transformed by advancements in open-source Large Language Models (LLMs) such as Llama2(Touvron et al., 2023), VicunaPeng et al. (2023), and MistralJiang et al. (2023). These potent models, undergirded by vast parameterizations and trained on varied data corpuses, have significantly broadened the scope of NLP applications.
Llama2, pioneered by Meta AI, provides a spectrum of options, with models ranging from 7 to 70 billion parameters. Renowned for its prowess in dialog-centric tasks, Llama2 leverages auto-regressive transformers fine-tuned with supervised learning and bolstered by reinforcement learning techniquesTouvron et al. (2023). Vicuna sets its mark as a specialized open-source LLM for chatbot applications, honed through datasets reflective of authentic conversational exchanges—flagships of platforms like ShareGPT. Vicuna’s clear training protocols and unique architecture, with models such as Vicuna-13B, tailor its performance to specific conversational needsPeng et al. (2023). In contrast, Mistral AI’s Mistral 7B stands as a versatile model catering to both English language processing and coding assignments. This model, with 7.3 billion parameters, outshines peers such as Llama 2 13B and Llama 1 34B across diverse benchmarks, thanks to its innovative grouped-query attention mechanism that optimizes the balance between resource economy and model reliabilityJiang et al. (2023).
Accompanying these is the "Chain of Thought" method Wei et al. (2022), another vanguard in language generation. Collectively, models like Llama2, Vicuna, Mistral, and Chain of Thought fortify a varied array of LLMs. Their contribution extends beyond improving NLP tasks; they are poised to pioneer new horizons in content generation and redefine our engagement with search technologies.
8 Conclusion
In this investigation, we initially incorporated LLMs into the realm of multi-intent SLU, thus tailoring the LM-MixATIS and LM-MixSNIPS datasets to effectively align with the proficiencies of comprehensive models. Concurrently, we introduced two unique model series: EN-LLM and ENSI-LLM. The results of our research indicate that LLMs yield competitive outcomes when juxtaposed with robust baseline models prevalent in this field. Upon further examination of the impact of data rate and intent quantity, we discovered certain limitations associated with LLMs. In addition to these findings, we formulated evaluation metrics, specifically ESA and CSA, which prove crucial in accurately assessing the performance of LLMs.
9 Limitations
(1) Impact of LoRA on Performance: While leveraging Low-Rank Adaptation (LoRA) affords improved training efficiency and reduced memory usage, it does not always achieve the full performance potential afforded by tuning all model parameters. LoRA strictly focuses on the most impactful parameters, which, though efficient, can sometimes lead to a loss of subtle linguistic detail that comprehensive parameter tuning might capture. (2) Prospects for Improvement through Data Curation and Prompt Optimization: Our current research framework does not extend to the advanced strategies of selective data curation or intricate prompt engineering. Recognizing this as a limitation, we propose that future investigations will embrace these crucial techniques.
Appendix A Appendix
Model Slot(F1) Intent(Acc) Overall(Acc)
EN-ChatGPT (5-shot) 64.0 54.1 14.6
EN-Vicuna-13B 87.9 83.6 50.8
EN-Llama-2-13B 87.9 81.0 49.5
EN-Mistral-8*7B 87.3 80.7 48.4
Table 5: Performance on the MixATIS dataset using various Large Language Models.
Model Slot(F1) Intent(Acc) Overall(Acc)
EN-ChatGPT(5-shot) 62.9 83.9 12.7
EN-Vicuna-13B 95.9 97.5 80.7
EN-Llama-2-13B 96.7 97.8 83.3
EN-Mistral-8*7B 96.1 97.5 80.9
Table 6: Performance on the MixSNIPS dataset using various Large Language Models.
Note1: We also extend our evaluation to examine ChatGPT’s proficiency in a 5-shot learning context, paralleling the tests conducted with the LLMs outlined above. A distinctive aspect of our methodology with ChatGPT involves introducing an explicit intent and slot list to direct the model’s learning, enhancing its capability to align with specified NLP tasks.
Note2: Performance analysis across different scales of Large Language Models (LLMs) yields a compelling insight: within the MixATIS dataset, the transition to more substantial 13B and 8*7B models does not uniformly translate into significant gains over their 7B counterparts. This finding challenges the notion that escalating model size inherently overcomes the challenges in discerning slots and intents within single-domain datasets. Conversely, the multi-domain MixSNIPS dataset paints a different picture, where enhanced model sizes seemingly marshal improved outcomes. Notably, the EN-Vicuna-13B model marks a 1.8% increase in overall accuracy relative to its 7B counterpart, the EN-Llama-2-13B model outperforms by an impressive 4.4%, and the EN-Mistral-8*7B model uplifts its performance by 1.1% in overall accuracy. These contrasts suggest that larger models may have a distinctive edge in more straightforward, multi-domain settings, chiefly attributed to their superior aptitude for capturing and interpreting the complex mosaic of patterns within such varied datasets.